# Grassroots Consensus

### v1

@@ 1 Introduction
Example: Grassroots Federation. A key application of digital social contracts mentioned in Table 1 is the grassroots federation of digital communities. Grassroots Federation [33] aims to address the egalitarian formation and the fair democratic governance of large-scale, decentralized, sovereign digital communities, the size of the EU, the US, existing social networks, and even humanity at large. A grassroots federation evolves via the grassroots formation of digital communities and their consensual federation. Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more. Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a small assembly elected by sortition [12, 33] among its members. It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community. Such a consensus protocol should satisfy all the requirements listed below.
@@
Requirements of Grassroots Consensus. Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency. Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued one at a time—the low-throughput scenario—finalize each quickly and efficiently. As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless. At the same time, the protocol should also work efficiently and with low-latency on smartphones in a high-throughput scenario. Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput. In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by a conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
@@
A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [28], Cryptocurrencies [30, 19]) are blocklace-based, it would be very beneficial and integrable to also have a (v) blocklace-based consensus protocol. This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [16] and its extensions and improvements [4, 14].
We note that a blocklace [3] is not “just a DAG” the same way a blockchain is not “just a sequence”. In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$. An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer of $b^{\prime}$. Thus, similarly to a blockchain, the blocklace is:
@@
However, unlike the blockchain, which is extended competitively, the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other. In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution. Namely the intention is that the parties to a social contract will run the contract, as well as its underlying consensus protocol, on their smartphones. The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult. In particular, opening a direct TCP connection between two phones is either impossible, or impossible without the help of a proxy server, and even if established, it has to be re-established whenever one of the phones changes their IP address. Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations. Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (vi) UDP-ready.
The final requirement for a protocol for grassroots platforms is, naturally, that the protocol be (iv) grassroots. Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$. Any permissioned consensus protocol satisfies that. (ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$. For this to happen in a permissioned consensus protocol, the set of agents $P$ need to be “reconfigured” to include agents in $P^{\prime}\setminus P$. The question of reconfiguration has been studied extensively for decades [18, 10, 2, 34]. While the original informal exposition of the problem in the Paxos paper was with respect to a self-governed parliament [18], subsequent works in the context of permissioned consensus assume that the initial ‘permissioned’ agents are determined by an external authority, and so is their ‘reconfiguration’. Proof-of-Stake protocols such as Ethereum 2.0 offer a stake-based approach to reconfigure the participating agents on every epoch [17, 6].
@@
Paper structure. Section 2 presents the Grassroots Consensus protocol assuming a reliable network. The protocol is presented assuming a prevailing constitution that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony. Section 3 proves the safety and liveness of Grassroots Consensus. Section 4 extends the protocol for eventually-reliable networks, making it UDP-ready. Section 5 proves that the Grassroots Consensus protocol is indeed grassroots. Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present a distinct democratic constitutional amendment process for each parameter. Section 7 concludes.
@@ 2.3 The Protocol
issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) a nack-block was received for $b^{\prime}$, or (iii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@ 5 Grassroots Consensus is Grassroots
Note that the two requirements cannot even be expressed in the context of the basic (non-constitutional) Grassroots Consensus protocol, as the population $P$ is given and fixed. And that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.

### v2

@@ 1 Introduction
Example: Grassroots Federation. A key application of digital social contracts mentioned in Table 1 is the grassroots federation of digital communities. Grassroots Federation [46] aims to address the egalitarian formation and the fair democratic governance of large-scale, decentralized, sovereign digital communities, the size of the EU, the US, existing social networks, and even humanity at large. A grassroots federation evolves via the grassroots formation of digital communities and their consensual federation. Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more. Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a similarly-small assembly elected by sortition [20, 46] among its members. It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community. Such a consensus protocol should satisfy all the requirements explained next.
@@
Requirements of Grassroots Consensus. Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency. Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued at low-throughput, finalize each quickly and efficiently. As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless.
At the same time, the protocol should also work efficiently and with low-latency on smartphones in a high-throughput scenario. Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput. In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by an overly-conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
@@
A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [41], Cryptocurrencies [43, 27]) are blocklace-based, it would be very beneficial and integrable to also have a (iv) blocklace-based consensus protocol. This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [24] and its extensions and improvements [5, 22]. We note that a blocklace [3] is not “just a DAG” the same way a blockchain is not “just a sequence”. In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$. An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer to $b^{\prime}$. Thus, similarly to a blockchain, the blocklace is:
@@
However, unlike the blockchain, which is extended competitively, with the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other. In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution. Namely, the intention is that the parties to a social contract will run both the contract and its underlying consensus protocol on their smartphones. The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult. In particular, opening a direct TCP connection between two phones is either impossible, or impossible without the help of a proxy server, and even if established, it has to be re-established whenever one of the phones changes their IP address. Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations. Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (v) UDP-ready.
The final sixth requirement for a protocol for grassroots platforms is, naturally, that the protocol be (vi) grassroots. Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$. Any permissioned consensus protocol satisfies that. (ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$.
Any platform that operates on a shared global resource or employs a global replicated (Blockchain [31]), or distributed (IPFS [7], DHT [35]) shared data structure, or distributed pub/sub systems with a global directory [13, 14, 8], are all not grassroots. Server-based federated systems such as Mastodon [34] are also not grassroots, informally due to the control exerted by each server-operator on its members; formally, since a group of people of any size cannot function without first connecting to such a server.
For a permissioned consensus protocol to be grassroots, the set of agents $P$ needs to be “reconfigured” to include agents in $P^{\prime}\setminus P$. The question of reconfiguration has been studied extensively for decades [26, 17, 2, 47]. While the original informal exposition of the problem in the Paxos paper was with respect to a self-governed parliament [26], subsequent works in the context of permissioned consensus assume that the initial ‘permissioned’ agents are determined by an external authority, and so is their ‘reconfiguration’. Proof-of-Stake protocols such as Ethereum 2.0 offer a stake-based approach to reconfigure the participating agents on every epoch [25, 10].
@@
Paper structure. Section 2 presents the Grassroots Consensus protocol assuming a reliable network, as well as a prevailing constitution $(P,\sigma,\Delta)$, that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony. Section 3 proves the safety and liveness of Grassroots Consensus. Section 4 extends the protocol for eventually-reliable networks, making it UDP-ready. Section 5 proves that the Grassroots Consensus protocol is indeed grassroots. Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present democratic processes geared for amending each of the three parameters. Section 7 concludes.
@@ 2.3 The Protocol
issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@ 5 Grassroots Consensus is Grassroots
Note that the two requirements cannot even be expressed in the context of the basic (non-constitutional) Grassroots Consensus protocol, as the population $P$ is given and fixed. Also note that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
@@ 6.2 Democratic Constitutional Amendment
Once $P$ can be amended, the question of sybils (fake and duplicate identities) entering $P$, in addition to Byzantines, comes to the fore. First, since one Byzantine agent, having penetrated $P$ with multiple identities, can cause as much harm as multiple Byzantine agents. In fact, the principal goal of permissionless protocols, both Proof-of-Work [32] and Proof-of-Stake [25, 10], is to augment consensus with mechanisms that make it sybil-resilient. Second, since a person controlling multiple digital identities in $P$—even if all correct at the consensus level—will have undue influence on the meta-level democratic processes, violating one person – one vote.
Existing large-scale digital communities suffer from extensive penetration of sybils. For example, Facebook removes around 1Bn fake accounts every quarter [48]. Most work on sybil detection and elimination employs top-down approaches suitable for global platforms [50, 18, 37, 12, 4, 6], but by their nature do not provide support for the democratic formation and conduct of digital communities. For grassroots consensus, egalitarian sybil-repelling protocols for the grassroots formation of digital communities [33, 39] may be more appropriate. They are based on mutual trust among the people forming a community and their willingness to backup this trust with mutual sureties. These protocols ensure, under certain conditions, that a digital community can grow indefinitely while retaining a bounded fraction of sybils.
A complementary line of work [38, 29], relevant to the meta-level democratic processes, shows how a community can employ sybil-resilient democratic governance: How the safety and liveness of digital democratic processes can be ensured despite a bounded penetration of sybils, even in the face of partial participation by its non-sybil members. Together, the two lines of work offer a comprehensive solution for (i) sybil-repellence during the egalitarian formation of digital communities and (ii) sybil-resilience in the democratic governance of digital communities, both needed during by the meta-level processes of constitutional amendment of Grassroots Consensus, as discussed next.